# Poincare field theory for massless particles

### v1

@@ Abstract:
Our main proposition is that field equations for all spins can be obtained from Casimir eigenvalue equations for Poincare group. We have already confirm that statement for massive scalar, spinor and vector fields in Ref.[1]. In the present article we are going to confirm this statement for massless vector, second and fourth rang tensor fields. In particular we will obtain Maxwell equations and Einstein equation in weak field approximation.
As is well known, Wigner define a particles as irreducible representation of Poincare group [2, 3]. But, as Weinberg noted in [4] irreducible representations for massless vector field with helicities ±$\pm 1$ do not exist. In the present article we will conform this statement for a wide class of massless fields. They are not Lorentz invariant since their Lorentz transformations have additional term in the form of gauge transformations. Such fields can appear in the theory only in the form which do not depend on corresponding gauge parameters. These forms are our equations of motion and they are by definition gauge invariant. So, the massless case is significantly different from massive one.
@@ 1 Introduction
In the previous article [1] we introduced equations of motion for massive particles with arbitrary spin as eigenvalue equations for Casimir operators of Poincare group. Here we will develop appropriate approach for massless particles. Although in massless case there is one Casimir operator (helicity), covariant form of its definition contains Pauli-Lubanski vector $W_{a}$ which has three independent components ($P^{a}W_{a}=0$). So, we have three equations and their compatibility produces significant differences from massive case.
@@
This feature has unexpected consequences discovered by Weinberg [4] that Lorentz transformations of some physically relevant fields (like vector and symmetric second rang tensor fields) have additional term in the form of gauge transformation. So, for these fields we can obtain actions as gauge invariant expressions of eigenvectors $\Psi^{A}_{i}$.
@@ 2.1 Representation for Pauli-Lubanski vector
Using representations for momenta and spin generators and the fact that they commute, we can find representation for Pauli-Lubanski vector.
@@ 2.2 Standard momentum $k^{a}$, space inversion and little group
It is useful to work in frame of standard momentum $k^{a}$. In massless case, when Lorentz invariant function of momentum vanish, $p^{2}=0$, we can chose $k^{a}=(1,0,0,-1)$. Then we can express any momentum $p^{a}$ as Lorentz transformation of $k^{a}$
@@
Since $P^{a}W_{a}=0$, Pauli-Lubanski vector has three components and for standard momentum they are
@@
Commutation relations of little group generators have a form
@@ 2.3 Principle field equations
We claim that this equations contains all massless free field equations with arbitrary helicity, known in classical field theory.
@@
Let us stress that all massless free field equations considered in classical field theory [4, 5, 6, 7, 8] can be obtained from Eqs.(2.3). In the present article we will confirm this statement for most important massless fields.
@@ 2.6 Solution of $(S_{12})^{A}{}_{B}$ equation
Next we can construct projection operators $(P_{i})^{A}{}_{B}$ on subspaces $\Xi_{i}$ (we will call it eigenspace) with dimension $d_{i}=dim\Xi_{i}$ corresponding to helicities $\lambda_{i}$
@@
Generally, (see section 8 for details) with the help of projection operators, for tensor fields of rang $n$ we can rewrite equation (20) in the form
@@
Note that for tensor fields of rang $n$ relation $\sum_{i=-n}^{n}d_{i}=d_{0}+2\sum_{i=1}^{n}d_{i}=4^{n}$ is satisfied and spectrum of helicities is
@@ 2.7.2 Gauge invariant fields
Next we will consider the most important physical cases of massless fields: scalar field, vector field with helicities $\lambda=\pm 1$ (photon) and symmetric second rank tensor with helicities $\lambda=\pm 2$ (metric). We will also consider second rank tensor with helicities $\lambda=\pm 1$ (field strength) and fourth rank tensor with helicities $\lambda=\pm 2$ (Riemann curvature).
@@ 4.1 Eigenvalues and projection operators
Note that eigenspace $\Xi_{0}$ corresponding to $\lambda_{0}=0$ is two dimensional $d_{0}=dim\Xi_{0}=2$ and consequently corresponding eigenvecor carries two degrees of freedom. The eigenspaces $\Xi_{\pm 1}$ corresponding to $\lambda_{\pm 1}=1$ are one dimensional $d_{\pm 1}=dim\Xi_{\pm 1}=1$ and consequently each of corresponding eigenvecors carries one degree of freedom.
@@ 4.2 Basic vectors
Note that tensors with highest helicity (for example tensor of rang $n$ with helicities $\lambda_{\pm n}=\pm n$) have a form $\breve{p}_{\pm}^{a_{1}}\breve{p}_{\pm}^{a_{2}}\cdots\breve{p}_{\pm}^{a_{n}}\mathbf{T}$. Consequently, tensors with highest helicity are one dimensional and they are symmetric in all indices. We will confirm that in all examples.
@@ 4.6 Action for massless vector field with helicities $\lambda=\pm 1$ produces Maxwell equations
5 Second rang tensor
In this section we will consider general features of second rang tensor.
@@ 5.1 Projection operators
In the case of second rang tensor we have $A\to(ab)$ and $\Psi^{A}\to T^{ab}$. We will keep notations $(P_{i})^{ab}{}_{cd}$ for projectors and $(S_{12})^{ab}{}_{cd}$ for spin operator of second rang tensor. Since we will use abbreviated notation we introduce new notations for vector operators, $(P_{0})^{a}{}_{b}=(\pi_{0})^{a}{}_{b}$ and $(P_{\pm 1})^{a}{}_{b}=(\pi_{\pm})^{a}{}_{b}$ for projectors and $(S_{12})^{a}{}_{b}=\rho^{a}{}_{b}$ for spin operator.
According to (2) in abbreviated notation representation of spin operator for second rang tensor takes a form
@@ 5.2 Solutions of $S_{12}$ equation
For second rang tensor characteristic matrix takes the form
@@
6 Second rang tensor with helicities $\lambda=\pm 2$ and general relativity in weak field approximation
Second rang tensor with helicities $\lambda=\pm 2$ is of particular importance since its symmetric part (metric) contribute to general relativity.
6.1 Gauge transformation for highest helicities $\lambda=\pm 2$ of second rang tensors
Each component of the second rang tensor with highest helicities $\lambda=\pm 2\,$
@@ 6.1 Gauge transformation for highest helicities $\lambda=\pm 2$ of second rang tensors
With the help of (2) we can express little group generators for massless second rang tensors $(\Pi_{1})^{ab}{}_{cd}$ and $(\Pi_{2})^{ab}{}_{cd}$ in terms of corresponding generators for vectors $(\Pi_{1})^{a}{}_{b}$ and $(\Pi_{2})^{a}{}_{b}$
@@
Here we came to a similar conclusion for highest helicities of second rang tensor as in the case of vector field: it is not physical since it is not Lorentz invariant.
@@ 6.2 General relativity in weak field approximation
As in the case of vector fields, components of massless second rang tensors of opposite helicities $\lambda=2$ and $\lambda=-2$ are related by transformation of space inversion. Since gravitation interaction is symmetric under space inversion it is common to treat both components of incomplete irreducible representation of proper Poincare group as a single particle called graviton. So, instead of two polarizations $T_{+2}^{ab}(k)$ and $T_{-2}^{ab}(k)$, both with one degree of freedom, we will introduce one symmetric tensor $h^{ba}(k)=h^{ab}(k)$ with two degrees of freedom, which is incomplete irreducible representation of Poincare group with space inversion. Using (11) we obtain
@@
7 Second rang tensor with helicities $\lambda=\pm 1$ produces Maxwell equations
Beside highes helicities $\lambda=\pm 2$, second rang tensor has also components with helicities $\lambda=\pm 1$. We expect that these components describe Maxwell equations in terms of field strength.
7.1 Second rang tensor with helicities $\lambda=\pm 1$
The components of second rang tensor with helicities $\lambda=\pm 1$ have a form
@@ 7.1 Second rang tensor with helicities $\lambda=\pm 1$
7.2 Gauge transformation for second rang tensor with helicities $\lambda=\pm 1$ and Maxwell equations
Using gauge transformations for basic vectors $\delta k^{a}=0$ and $\delta\breve{p}_{\pm}^{a}=k^{a}\varepsilon_{\pm}$ we can obtain gauge transformations for symmetric and antisymmetric parts of second rang tensor with helicities $\lambda=\pm 1$
@@ 7.2.2 Antisymmetric tensor with helicities $\lambda=\pm 1$ and Maxwell equations
8 Fourth rang tensor with helicities $\lambda=\pm 2$ produces general relativity in weak field approximation
As symmetric second rang tensor with helicities $\lambda=\pm 1$ describe electromagnetic interaction through field strength, we can expect that fourth rang tensor with helicities $\lambda=\pm 2$ and appropriate symmetries describe gravitation interaction through Riemann tensor. In the case of fourth rang tensor we have $A\to(abcd)$ and $\Psi^{A}\to T^{abcd}$. In this section to simplify notation we will keep index $A$ instead of four indices $abcd$. When we include symmetry properties we will go back to the notation with indices $abcd$.
8.1 Fourth rang tensor
@@
where $T^{A}$ is fourth rang tensor, $T^{A}=T^{abcd}$. The problem comes down to that of vector fields. According to (2) in abbreviated notation spin operator takes the form
@@
which produces spectrum of helicities for four rang tensor
$\lambda_{0}=0\,,\qquad\lambda_{\pm i}=\pm i\,.\qquad i=(1,2,\cdots,4)$ (10)
The eigenfunctions, which are our candidates for irreducible representations, for four rang massless tensors are
$T_{i}^{A}(k)=(P_{i})^{A}{}_{B}T^{B}(k)\,.\qquad(i=0,\pm 1,\pm 2,\cdots,\pm 4)$ (11)
Note that approach of this subsection can be easily generalized to tensors of arbitrary rang.
@@ 8.2 General relativity in terms of Riemann tensor
In this subsection we will find irreducible representations of fourth rang tensor with helicity $\lambda=\pm 2$. First we obtain irreducible representations of proper Poincare group, $T_{\pm 2(-)}^{abcd}(k)$ and then of wider group which combined Poincare group and parity, $R^{abcd}(k)$. The last one is linearized Riemann tensor. Finally, we obtain Einstein equations in vacuum.
@@ 8.2.1 Irreducible representations of proper Poincare group
The fourth rang tensor with helicities $\lambda=\pm 2$ has a form
@@
As well as in the case of second rang tensor with helicities $\lambda=\pm 1$, to describe particles with helicities $\lambda=\pm 2$ with the help of fourth rang tensor, in every place we will substitute $\pi_{0}$ with its gauge invariant part $\pi_{0-}$. Then fourth rang tensor with helicities $\lambda=\pm 2$ has two terms: one with $2$ vectors $\breve{p}_{\pm}^{a}$ and $2$ vectors $k^{a}$ and the other with $3$ vectors $\breve{p}_{\pm}^{a}$ and one vector $\breve{p}_{\mp}^{a}$
@@
This tensor is symmetric under exchange of the two pairs of indices ($ab$) and ($cd$). We can rewrite it in terms of second rang symmetric tensor $h_{\pm}^{ab}=\mathbf{T}_{\mp 4}^{R}\breve{p}_{\pm}^{a}\breve{p}_{\pm}^{b}$ as
@@ 8.2.4 Dimensions of the representations and degrees of freedom
Let us count dimensions of these representations. As we have already shown dimension of each of representation $T_{\pm 2}^{abcd}(k)$ defined in Eq.(8.2.1) is $d_{2}=28$, where we used the fact that subspace corresponding to $\pi_{0}$ is two dimensional. The first term is $6\cdot 2\cdot 2=24$ dimensional where $6$ is number of combination of $4$ things taken $2$ at a time and $2\cdot 2$ is dimension of subspace corresponding to two operators $\pi_{0}$. The second term is $4$ dimensional so that $24+4=28$. After substitution two dimensional projector $\pi_{0}$ with one dimensional $\pi_{0-}$ we obtain that $T_{\pm 2(-)}^{abcd}(k)$, defined in Eq.(13) is $6\cdot 1\cdot 1+4=10$ dimensional representation, which could be expected for symmetric second rang tensor.

### v2

@@ Abstract:
Our main proposition is that field equations for all spins can be obtained from Casimir eigenvalue equations for Poincare group. We have already confirm that statement for massive scalar, spinor and vector fields in Ref.[1]. In the present article we are going to confirm this statement for massless vector, second and fourth rank tensor fields. In particular we will obtain Maxwell equations and Einstein equation in weak field approximation.
As is well known, Wigner defined particles as irreducible representation of Poincare group [2, 3]. But, as Weinberg noted in [4] irreducible representations for massless vector field with helicities ±$\pm 1$ do not exist. In the present article we will conform this statement for a wide class of massless fields. They are not Lorentz invariant since their Lorentz transformations have additional term in the form of gauge transformations. Such fields can appear in the theory only in the form which do not depend on corresponding gauge parameters. These forms are our equations of motion and they are by definition gauge invariant. So, the massless case is significantly different from massive one. In the end we will show that our approach can reproduce main contributions from well known articles with highest helicity.
@@ 1 Introduction
In the previous article [1] we introduced equations of motion for massive particles with arbitrary spin as eigenvalue equations for Casimir operators of Poincare group. Here we will develop appropriate approach for massless particles. Although in massless case there is one Casimir operator (helicity), covariant form of its definition contains Pauli-Lubanski vector $W_{a}$ which has three independent components (since $P^{a}W_{a}=0$). So, we have three equations and their compatibility produces significant differences from massive case.
@@
This feature has unexpected consequences discovered by Weinberg [4] that Lorentz transformations of some physically relevant fields (like vector and symmetric second ranktensor fields) have additional term in the form of gauge transformation. So, for these fields we can obtain actions as gauge invariant expressions of eigenvectors $\Psi^{A}_{i}$.
@@
The main idea of this work is to show that principle field equations produce all known field equations. So, we will first obtain standard equations for helicities $1$ and $2$. There is also significant interest in massless fields with higher spin. The theory of particles with highest spin was developed by Fierz and Pauli [5]. In Ref. [2] and [3] it has been shown that the one-particle states is an irreducible, unitary representation of the Poincare group. The set of a minimum auxiliary traceless, symmetric tensor fields was proposed in Ref.[6]. Since it is convenient to have an action for a field theory in order to perform quantization, equations of motion and action was found in Ref.[7]. There is also interesting approach based on a hierarchy of generalized Christoffel symbols by de Wit and Freedman [8]. In all of the above articles, fields under consideration satisfy Fierz-Pauli constraints [5] which are the basis of these articles. It means that field is totally symmetric tensor of rank $n$, which is traceless, divergence free and satisfying the Klein-Gordon equation.
In the present article we offer more fundamental approach. For arbitrary massless tensor we developed procedure to find all possible helicities and structure od tensor for each given helicity. As a particular case, for given tensor of rank $n$ highest helicity is totally symmetric, traceless, divergence free and satisfying the Klein-Gordon equation. So, our highest helicity satisfy Fierz-Pauli constraints [5]. We also considered all non-highest helicity and offer theirs physical interpretation, see sections 7. and 8.
@@ 2.1 Representation for Pauli-Lubanski vector
Using representations for momenta and for spin generators as well as the fact that they commute, we can find representation for Pauli-Lubanski vector.
@@ 2.2 Standard momentum $k^{a}$, space inversion and little group
It is useful to work in frame of standard momentum $k^{a}$. In massless case, when Lorentz invariant function of momentum vanish, $p^{2}=0$, we can chose standard momentum in the form $k^{a}=(1,0,0,-1)$. Then we can express any momentum $p^{a}$ as Lorentz transformation of $k^{a}$
@@
Since $P^{a}W_{a}=0$, Pauli-Lubanski vector has three independent components and for standard momentum they are
@@
Commutation relations of little group generators have the form
@@ 2.3 Principle field equations
We claim that all massless free field equations with arbitrary helicity, considered in classical field theory [4, 9, 10, 11, 12] can be obtained from this equation. In the present article we will confirm this statement for most important massless fields.
@@ 2.6 Solution of $(S_{12})^{A}{}_{B}$ equation
Next, we can construct projection operators $(P_{i})^{A}{}_{B}$ on subspaces $\Xi_{i}$ (we will call it eigenspace) with dimension $d_{i}=dim\Xi_{i}$ corresponding to helicities $\lambda_{i}$
@@
Generally, (see section 9 for details) with the help of projection operators, for tensor fields of rank $n$ we can rewrite equation (20) in the form
@@
Note that for tensor fields of rank $n$ relation $\sum_{i=-n}^{n}d_{i}=d_{0}+2\sum_{i=1}^{n}d_{i}=4^{n}$ is satisfied and spectrum of helicities is
@@ 2.7.2 Gauge invariant fields
Next we will consider the most important physical cases of massless fields: scalar field, vector field with helicities $\lambda=\pm 1$ (photon) and symmetric second rank tensor with helicities $\lambda=\pm 2$ (metric). We will also consider second rank tensor with helicities $\lambda=\pm 1$ (field strength) and fourth rank tensor with helicities $\lambda=\pm 2$ (Riemann curvature). We will also comment gauge fields with higher helicities.
@@ 4.1 Eigenvalues and projection operators
Note that eigenspace $\Xi_{0}$ corresponding to $\lambda_{0}=0$ is two dimensional $d_{0}=dim\Xi_{0}=2$ and consequently corresponding eigenvector carries two degrees of freedom. The eigenspaces $\Xi_{\pm 1}$ corresponding to $\lambda_{\pm 1}=1$ are one dimensional $d_{\pm 1}=dim\Xi_{\pm 1}=1$ and consequently each of corresponding eigenvectors carries one degree of freedom.
@@ 4.2 Basic vectors
[…]
@@ 4.6 Action for massless vector field with helicities $\lambda=\pm 1$ produces Maxwell equations
5 Second rank tensor
In this section we will consider general features of second rank tensor.
@@ 5.1 Projection operators
In the case of second rank tensor we have $A\to(ab)$ and $\Psi^{A}\to T^{ab}$. We will keep notations $(P_{i})^{ab}{}_{cd}$ for projectors and $(S_{12})^{ab}{}_{cd}$ for spin operator of second ranktensor. Since we will use abbreviated notation we introduce new notations for vector operators, $(P_{0})^{a}{}_{b}=(\pi_{0})^{a}{}_{b}$ and $(P_{\pm 1})^{a}{}_{b}=(\pi_{\pm})^{a}{}_{b}$ for projectors and $(S_{12})^{a}{}_{b}=\rho^{a}{}_{b}$ for spin operator.
According to (2) in abbreviated notation representation of spin operator for second rank tensor takes a form
@@ 5.2 Solutions of $S_{12}$ equation
For second rank tensor characteristic matrix takes the form
@@
6 Second rank tensor with helicities $\lambda=\pm 2$ and general relativity in weak field approximation
Second rank tensor with helicities $\lambda=\pm 2$ is of particular importance since its symmetric part (metric) contribute to general relativity.
6.1 Gauge transformation for highest helicities $\lambda=\pm 2$ of second rank tensors
Each component of the second rank tensor with highest helicities $\lambda=\pm 2\,$
@@ 6.1 Gauge transformation for highest helicities $\lambda=\pm 2$ of second rank tensors
With the help of (2) we can express little group generators for massless second rank tensors $(\Pi_{1})^{ab}{}_{cd}$ and $(\Pi_{2})^{ab}{}_{cd}$ in terms of corresponding generators for vectors $(\Pi_{1})^{a}{}_{b}$ and $(\Pi_{2})^{a}{}_{b}$
@@
Here we came to a similar conclusion for highest helicities of second rank tensor as in the case of vector field: it is not physical since it is not Lorentz invariant.
@@ 6.2 General relativity in weak field approximation
As in the case of vector fields, components of massless second rank tensors of opposite helicities $\lambda=2$ and $\lambda=-2$ are related by transformation of space inversion. Since gravitation interaction is symmetric under space inversion it is common to treat both components of incomplete irreducible representation of proper Poincare group as a single particle called graviton. So, instead of two polarizations $T_{+2}^{ab}(k)$ and $T_{-2}^{ab}(k)$, both with one degree of freedom, we will introduce one symmetric tensor $h^{ba}(k)=h^{ab}(k)$ with two degrees of freedom, which is incomplete irreducible representation of Poincare group with space inversion. Using (11) we obtain
@@
7 Second rank tensor with helicities $\lambda=\pm 1$ produces Maxwell equations
Beside highes helicities $\lambda=\pm 2$, second rank tensor has also components with helicities $\lambda=\pm 1$. We expect that these components describe Maxwell equations in terms of field strength.
7.1 Second rank tensor with helicities $\lambda=\pm 1$
The components of second rank tensor with helicities $\lambda=\pm 1$ have a form
@@ 7.1 Second rank tensor with helicities $\lambda=\pm 1$
7.2 Gauge transformation for second rank tensor with helicities $\lambda=\pm 1$ and Maxwell equations
Using gauge transformations for basic vectors $\delta k^{a}=0$ and $\delta\breve{p}_{\pm}^{a}=k^{a}\varepsilon_{\pm}$ we can obtain gauge transformations for symmetric and antisymmetric parts of second rank tensor with helicities $\lambda=\pm 1$
@@ 7.2.2 Antisymmetric tensor with helicities $\lambda=\pm 1$ and Maxwell equations
8 Fourth rank tensor with helicities $\lambda=\pm 2$ produces general relativity in weak field approximation
As symmetric second rank tensor with helicities $\lambda=\pm 1$ describe electromagnetic interaction through field strength, we can expect that fourth rank tensor with helicities $\lambda=\pm 2$ and appropriate symmetries describe gravitation interaction through Riemann tensor. In the case of fourth rank tensor we have $A\to(abcd)$ and $\Psi^{A}\to T^{abcd}$. In this section to simplify notation we will keep index $A$ instead of four indices $abcd$. When we include symmetry properties we will go back to the notation with indices $abcd$.
8.1 Fourth rank tensor
@@
where $T^{A}$ is fourth rank tensor, $T^{A}=T^{abcd}$. The problem comes down to that of vector fields. According to (2) in abbreviated notation spin operator takes the form
@@
which produces spectrum of helicities for four rank tensor
$\lambda_{0}=0\,,\qquad\lambda_{\pm i}=\pm i\,.\qquad i=(1,2,3,4)$ (10)
The eigenfunctions, which are our candidates for irreducible representations, for four rank massless tensors are
$T_{i}^{A}(k)=(P_{i})^{A}{}_{B}T^{B}(k)\,.\qquad(i=0,\pm 1,\pm 2,\pm 3,\pm 4)$ (11)
@@ 8.2 General relativity in terms of Riemann tensor
In this subsection we will find irreducible representations of fourth rank tensor with helicity $\lambda=\pm 2$. First we obtain irreducible representations of proper Poincare group, $T_{\pm 2(-)}^{abcd}(k)$ and then of wider group which combined Poincare group and parity, $R^{abcd}(k)$. The last one is linearized Riemann tensor. Finally, we obtain Einstein equations in vacuum.
@@ 8.2.1 Irreducible representations of proper Poincare group
The fourth rank tensor with helicities $\lambda=\pm 2$ has a form
@@
As well as in the case of second rank tensor with helicities $\lambda=\pm 1$, to describe particles with helicities $\lambda=\pm 2$ with the help of fourth rank tensor, in every place we will substitute $\pi_{0}$ with its gauge invariant part $\pi_{0-}$. Then fourth rank tensor with helicities $\lambda=\pm 2$ has two terms: one with $2$ vectors $\breve{p}_{\pm}^{a}$ and $2$ vectors $k^{a}$ and the other with $3$ vectors $\breve{p}_{\pm}^{a}$ and one vector $\breve{p}_{\mp}^{a}$
@@
This tensor is symmetric under exchange of the two pairs of indices ($ab$) and ($cd$). We can rewrite it in terms of second rank symmetric tensor $h_{\pm}^{ab}=\mathbf{T}_{\mp 4}^{R}\breve{p}_{\pm}^{a}\breve{p}_{\pm}^{b}$ as
@@ 8.2.4 Dimensions of the representations and degrees of freedom
Let us count dimensions of these representations. As we have already shown dimension of each of representation $T_{\pm 2}^{abcd}(k)$ defined in Eq.(8.2.1) is $d_{2}=28$, where we used the fact that subspace corresponding to $\pi_{0}$ is two dimensional. The first term is $6\cdot 2\cdot 2=24$ dimensional where $6$ is number of combination of $4$ things taken $2$ at a time and $2\cdot 2$ is dimension of subspace corresponding to two operators $\pi_{0}$. The second term is $4$ dimensional so that $24+4=28$. After substitution two dimensional projector $\pi_{0}$ with one dimensional $\pi_{0-}$ we obtain that $T_{\pm 2(-)}^{abcd}(k)$, defined in Eq.(13) is $6\cdot 1\cdot 1+4=10$ dimensional representation, which could be expected for symmetric second ranktensor.